OpenStudy (anonymous):

The partial fraction decomposition of a rational function is -3/(x+1) + (-2x) / (x^2+9). Evaluate the indefinite integral

OpenStudy (anonymous):

I did the evaluation and found the answer to be -log(x^2+9)-3log(x+1) but webwork says it's wrong

OpenStudy (irishboy123):

if you mean \(\large \int \frac{-3}{(x+1)} + \frac{-2x}{(x^2+9)} \ dx \) looks right, if you added an integration constant maybe you need to manipulate a bit to please the software?

OpenStudy (anonymous):

Well the software returns an extra set of parentheses, making the answer\[-(\log(x^2+9))-3\log(x+1)\] It seems like it should be right, but it just won't accept it

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

hi @RunawayGalaxy or you can write \(ln \frac{1}{x^2 + 9} + ln \frac {1}{(x+1)^3} + C\) dunno, looks like a glitch, sure a human would be very happy with your work, webwork is presumably some kinda online thing....

OpenStudy (anonymous):

Yeah, it's the online homework program for my Calc class. Well thanks for the help, I'll try that, and email my professor if it doesn't work!

OpenStudy (irishboy123):

well, i stuffed it into Wolfram to eliminate "completely losing the plot" risk... and the Wolf agrees with you! http://www.wolframalpha.com/input/?i=+%5Cint+%5Cfrac%7B-3%7D%7B%28x%2B1%29%7D+%2B+%5Cfrac%7B-2x%7D%7B%28x%5E2%2B9%29%7D+++dx